OpenStudy (ksaimouli):

find the domain of the function f.use limits to describe the behaviour of f at valus of x not in its domain 1/x+3

OpenStudy (ksaimouli):

plz explain me

OpenStudy (slaaibak):

For what value of x will you divide by 0?

OpenStudy (ksaimouli):

?

OpenStudy (ksaimouli):

\[1\div x+1\]

OpenStudy (ksaimouli):



OpenStudy (slaaibak):

Cool. \[f(x) = {1 \over x+3}\] Where do you think will this function have a discontinuity? Aka, for what values of x is this function undefined?

OpenStudy (ksaimouli):

-3

OpenStudy (slaaibak):

Yes. So, do you agree this function is defined for every value of x, except for -3?

OpenStudy (ksaimouli):

yup

OpenStudy (ksaimouli):

so ow to write end bhaviour

OpenStudy (ksaimouli):

i am stuck with that

OpenStudy (slaaibak):

Then we can agree that the domain of the function is: \[(-\infty,-3) \cup (-3, \infty)\]

OpenStudy (slaaibak):

That means x goes from -infinity to JUST before three, meaning every value smaller than -3. and then it unites with every value larger than -3 to infinity

OpenStudy (ksaimouli):

thx slaaibak

OpenStudy (ksaimouli):

thank u ver much

OpenStudy (slaaibak):

For the limits: \[\lim_{x \rightarrow -3^-} { 1 \over x + 3} = -\infty\] similarly \[\lim_{x \rightarrow -3^+} { 1 \over x + 3} = \infty\] and lastly \[\lim_{x \rightarrow -3} { 1 \over x + 3} = Does NotExist\]

OpenStudy (ksaimouli):

ok got u but what happens in this case